# Is Your Imitation Learning Policy Better than Mine? Policy Comparison with Near-Optimal Stopping

### v1

@@ Abstract
Imitation learning has enabled robots to perform complex, long-horizon tasks in challenging dexterous manipulation settings. As new methods are developed, they must be rigorously evaluated and compared against corresponding baselines through repeated evaluation trials. However, policy comparison is fundamentally constrained by a small feasible sample size (e.g., 10 or 50) due to significant human effort and limited inference throughput of policies. This paper proposes a novel statistical framework for rigorously comparing two policies in the small sample size regime. Prior work in statistical policy comparison relies on batch testing, which requires a fixed, pre-determined number of trials and lacks flexibility in adapting the sample size to the observed evaluation data. Furthermore, extending the test with additional trials risks inducing inadvertent p-hacking, undermining statistical assurances. In contrast, our proposed statistical test is sequential, allowing researchers to decide whether or not to run more trials based on intermediate results. This adaptively tailors the number of trials to the difficulty of the underlying comparison, saving significant time and effort without sacrificing probabilistic correctness. Extensive numerical simulation and real-world robot manipulation experiments show that our test achieves near-optimal stopping, letting researchers stop evaluation and make a decision in a near-minimal number of trials. Specifically, it reduces the number of evaluation trials by up to 40% as compared to state-of-the-art baselines, while preserving the probabilistic correctness and statistical power of the comparison. Moreover, our method is strongest in the most challenging comparison instances (requiring the most evaluation trials); in a multi-task comparison scenario, we save the evaluator more than 200 simulation rollouts.
@@ I Introduction
A particularly important aspect of policy evaluation is policy comparison, where two policies are repeatedly deployed in an environment to assess their relative performance. Policy comparison forms the foundation of robot learning as “an empirical science” [28], enabling researchers to objectively measure the scientific progress of the field. Nevertheless, this setting introduces an additional source of stochasticity due to random outcomes of the second policy, making the reliability of the comparison more challenging to ensure.
@@
To address these challenges, we propose a novel sequential testing framework named STEP (Sequential Testing for Efficient Policy Comparison) for rigorously comparing performance of imitation learning policies. Unlike batch testing, our approach allows the number of trials to be varied within a given experimental budget. This critical feature offers two practical advantages. First, the evaluator can stop conducting trials early without sacrificing the probabilistic correctness of the comparison if enough statistical evidence is accumulated quickly in favor of $\pi_{1}$ (or $\pi_{0}$). Second, it reduces the epistemic risk of overconfident (and potentially incorrect) conclusions when $\pi_{0}$ and $\pi_{1}$ are closely matched, since the test will abstain from making a decision if statistical evidence remains low after all the planned trials have been executed. Alternatively, the evaluator may safely append additional trials to the original samples and re-conduct statistical analysis on all the results obtained thus far without inadvertent p-hacking. We demonstrate these advantages through simulation and real-world robot manipulation experiments. Furthermore, extensive numerical experiments show that our STEP significantly outperforms state-of-the-art (SOTA) sequential methods, reducing the required number of trials by up to 40% without sacrificing probabilistic correctness. The specific contributions of this paper are as follows:
  • We propose STEP, a novel sequential statistical framework for evaluating relative performance of two policies with tunable probabilistic correctness.
@@ II Preliminaries
Null Hypothesis $\begin{split}\text{Null Hypothesis }\mathbb{H}_{0}:&\hskip 5.69054ptp_{1}~{}\leq~{}p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{0}\\ \text{Alternative Hypothesis }\mathbb{H}_{1}:&\hskip 5.69054ptp_{1}>p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{1}.\end{split}$ (3)
@@
Finally, while the preceding development has focused on one-sided testing, the comparison problem naturally admits a bidirectional version allowing decisions for membership in either the alternative set (RejectNull) or the null set (AcceptNull). Note that the decision AcceptNull formally amounts to rejecting the null of a “flipped” version of Eq. 3. In all subsequent discussion we will implicitly utilize the bidirectional test, which will allow for the decision
@@ IV Methodology
There are several important practical considerations to constructing a near-optimal solution to Eq. 4. We present the concrete challenges first, and then discuss the technical innovations that account for them. Throughout, we will leverage significant mathematical structure in the testing problem. Where insightful or intuitive, this will be explained in situ. See Section IX-E in the Supplement for additional details.
@@ IV-C Type-I Error Control
The composite nature of the null hypothesis (which contains all pairs $\mathcal{H}_{0}=\{(p_{0},p_{1})\in[0,1]^{2}\text{ }\rvert\text{ }p_{0}~{}\geq~{}p_{1}\}$) means that the decision-making problem can be thought of (informally) as analogous to distributionally robust control, where the uncertainty is over the particular worst-case $(p_{0},p_{1})\in\mathcal{H}_{0}$. Controlling the Type-I Error in policy comparison then amounts to controlling the Type-I Error uniformly (robustly) for all $h\in\mathcal{H}_{0}$. Suppose that some rejection region for the first $t-1$ steps has been obtained with accumulated risk $\alpha_{n-1}$, and we are interested in bounding the Type-I Error for $n$ by some $\alpha_{n}>\alpha_{n-1}$. Given the notion of stopping and continuation regions introduced in the previous section, this is expressed mathematically as:
@@
The dependence on the (probabilistic) rejection region for $t-1$ is made explicit in Eq. 7, reflecting the internal dynamic structure. For example, if some state $\mathbf{x}_{n-1}=(a,b,n-1)$ is rejected at $n-1$ with a non-zero probability, then the 1-step reachable states (e.g., $(a+1,b,n)$) under the dynamics Eq. 5 are less likely to be feasible at $n$. The presence of many (infinite) elements $h$ in the set of null hypotheses $\mathcal{H}_{0}$ makes verification of Type-I Error control challenging a priori, even for a single particular $n$. However: monotonicity, symmetry, and smoothness properties of this family of testing problems (noted in, for example, [13]) allow for efficient discretization procedures that preserve safety. Specifically, it suffices to consider a discrete set of the “worst-case” nulls ^$\hat{\mathcal{H}}_{0}=\{(p^{(1)},p^{(1)}),(p^{(2)},p^{(2)}),\cdots,(p^{(M)},p^{(M)})\}$, where $10~{}\leq~{}p^{(1)}\cdots~{}\leq~{}p^{(M)}~{}\leq~{}$ 1. Further details of this discretization process is given in Section IX-E2 in the Supplement. Thus, Eq. 7 reduces to a finite set of individual inequality constraints.
@@ IV-E Tractable Optimization
The optimization problem Eq. 10 is a simple linear program that can be efficiently solved by a standard optimization software; it is a maximization of $\lVert\mathbf{w}_{n}\rVert_{1}$ over the nonnegative orthant, subject to additional linear inequality constraints to control Type-I Error. Because $P_{n}$ depends on the rejection regions of $n-1$, the optimization needs to be performed sequentially for each $n$ in increasing order. Nevertheless, all the computation can be performed offline prior to running the actual policy evaluation. The fact that the optimization problems are sequentially solvable is owed to the isolation of the risk accumulation rate $f(n)$ as a tunable parameter; otherwise, the rejection regions would be generally coupled across $n$ and the optimization would be more challenging.
@@ V-A Baseline Procedures
The baselines constitute the SOTA sequential analysis methods described in Lai [30] and Lai and Zhang [31] (termed “Lai”) and the Safe, Anytime-Valid Inference (SAVI) method of Turner and Grünwald [50], which is specifically tailored to contingency table (i.e., policy comparison) problems (termed “SAVI”). Similar to STEP, Lai is a valid sequential method under a pre-determined $N_{\max}$ and is asymptotically optimal as $N_{\max}$ tends to infinity. SAVI does not require a fixed $N_{\max}$ a priori and still maintains validity for an arbitrary large $N_{\max}$. Both methods are widely applicable to practical testing problems. Additionally, an Oracle Sequential Probability Ratio Test (SPRT) [55] is run using the true singleton alternative and associated worst-case singleton null; this method represents a near-optimal procedure for easier problems where feasible methods quickly approach a terminal power close to 1 and still serves as a reasonable benchmark in harder problems. We emphasize that in practical cases, the Oracle method is infeasible to the evaluator; it is included to give a conservative estimate of the optimality gap of each method. Additional information about each baseline is included in Section VI.
@@ V-B Numerical Simulation
The algorithms are first validated on multiple pairs of $(N_{\text{max}},\alpha^{*})$. The ensuing numerical results will use the case $N_{\text{max}}=500,\alpha^{*}=0.05$; similar figures for $N_{\text{max}}=100,\alpha^{*}=0.05$ are included in Section IX-B in the Supplement. For each method, we compute for each of the 45 alternatives the following characteristics: (1) Type-I Error – the fraction of associated worst-case null trajectories which have incorrectly rejected the null; (2) Terminal Power – the fraction of alternative trajectories which have correctly rejected the null by step $N_{\text{max}}$; (3) Cumulative Power – a visualization of the fraction of alternative trajectories which have correctly rejected the null by step $t$ for all $t\in\{0,1,...,N_{\text{max}}\}$. Note that another natural evaluation metric, the expected time-to-decision ($\mathbb{E}[n_{\text{stop}}]$), can be derived from the cumulative power as the area between the curve and the constant $y=1$.
@@ V-B2 What are the unique advantages of STEP in comparison problems of varying difficulty?
We now demonstrate the downside of the Lai baseline procedure as compared to our method through the cumulative power of the procedure. In Fig. 5 (right), we illustrate the cumulative power on a hard evaluation case. This represents a second view of the observation presented in the terminal power analysis: the Lai procedure and our STEP each significantly outperform the SAVI procedure in the small-gap regime. Conversely, in Fig. 5 (center), we illustrate a difficult setting in which the Lai procedure struggles. In this setting the performance gap is again 10 percentage points, but the distribution is lower-variance and skewed as compared to that of Fig. 5 (right); as such, the Lai procedure cannot effectively adapt and our STEP significantly outperforms in terms of deciding more quickly for the alternative.
@@ V-C How sample efficient is STEP in practical policy comparison settings?
In addition to the numerical validation, we evaluate STEP through two sets of real-world robot evaluation experiments and a simulation experiment. In the first set of experiments, we compared policies with noticeable performance gaps to demonstrate the early-stopping capability of our approach. In the second set, we compared closely-competing policies to characterize necessary sample sizes for statistical validity when policy performance gaps become small. In the simulation experiment, we perform multi-task comparison of two SOTA imitation learning policies. See Section IX-A in the Supplement for more details on the hardware experiments.
@@ V-C1 Hardware Evaluation in the Large/Medium-Gap Regime
In FoldRedTowel, we compare two policy checkpoints from a single training run. The baseline policy $\pi_{0}$ was trained for 10000 gradient steps with an AdamW [34] optimizer and achieved the action mean-squared error (MSE) of 1.61E-3 on the validation set. The other policy $\pi_{1}$ continued the training for 70000 additional steps, yielding the validation action MSE of 1.35E-3. To evaluate each policy, five in-distribution (ID) initial conditions were chosen and repeated 10 times each, constituting 50 total trials. ID initial conditions visually resemble the ones that appear in the training dataset. As shown in Table I (rows 1-2), the empirical gap in success rates was 36 percentage points (56% to 92% success), suggesting that $\pi_{0}$ was under-trained. Each sequential method detected a significant difference at level $\alpha^{*}=0.05$ in 20 – 23 trials (apiece). That is, had the algorithms been active during collection, the last 27 – 30 rollouts per policy were unnecessary for confirming the improvement of the later checkpoint. Additionally, note that the Lai and our STEP sequential procedures were each tuned for an $N_{\text{max}}$ of 200 (row 1) and 500 (row 2) rollouts, meaning that if the gap was smaller, additional rollouts could have been run up to 200 or 500 per policy without compromising the validity of the decision.
@@
FoldRedTowel $\pi_{i}$ 0.05 500 50 0.560 0.920 20 20 23 17
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 200 50 0.400 0.820 13 35 14 12
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 500 50 0.400 0.820 13 35 18 12
@@
TABLE I: Empirical time-to-correct-decision for all hardware (top) and simulation (bottom) policy comparisons. The comparison type is described first; $\pi_{i}$ is comparing two policies, $\mathcal{D}^{i}_{s_{0},o_{0}}$ compares one policy under possible distribution shift. The utilized Type-I Error $\alpha^{*}$ and $N_{\text{max}}$ describe the constraints applied a priori by the evaluator (we underline to emphasize the change in $N_{\text{max}}$ for rows 1-2 and 3-4; observe that the sensitivity of the stopping times is very small). $N~{}\leq~{}N_{\text{max}}$ represents the amount of data available for the statistical analysis. We report the terminal empirical success rates (after $N$ trials) of each policy in each setting under $\hat{p}_{i}$ (this information is not available to any feasible algorithm). We do not have truth labels on this data; however, in all cases, every method arrived at the same decision, including the Oracle SPRT which has a priori access to privileged information $(\hat{p}_{0},\hat{p}_{1})_{N}$; this decision was Reject Null for all rows except the CarrotOnPlate tasks, which each returned Fail To Decide. We report the stopping times of all methods on the right of the table for every context; in all cases: lower is better. We put in bold any feasible method result that is near-optimal within ten trials (absolute) or 25% (relative) of the SPRT Oracle, which is not implementable by an evaluator. In the Multitask setting, we test $p_{1}>p_{0}$ uniformly across the preceding three tasks. This stopping time is the sum by column of the stopping times for the three tasks. Our method saves the evaluator over 200 trials in uniform certification over these three tasks as compared to either feasible baseline.
In CleanUpSpill, we compare the same policy on two different sets of initial conditions. For this task we reuse the evaluation results originally presented in [anonymous citation], which compare a set of ID initial conditions against the out-of-distribution (OOD) initial conditions. The ID set includes ten initial conditions with a white towel whereas the OOD set uses ten with a green towel. As shown in Table I (row 3), the empirical gap of 42 percentage points was detected in under 20 trials by all methods except the Lai procedure, though they were tuned (where applicable) to an $N_{\text{max}}$ nearly ten to twenty times larger; this demonstrates the significant reduction in sensitivity (from an evaluator’s standpoint) arising from setting $N_{\text{max}}$ versus choosing a batch size $N$. Furthermore, STEP’s efficiency only minimally degrades when $N_{\text{max}}$ is increased from 200 (row3) to 500 (row 4). In this setting, the best-performing sequential methods would have prevented the need for nearly 60 of the 100 total rollouts (30 of the 50 batch trials per distribution).

### v2

@@ Abstract
Imitation learning has enabled robots to perform complex, long-horizon tasks in challenging dexterous manipulation settings. As new methods are developed, they must be rigorously evaluated and compared against corresponding baselines through repeated evaluation trials. However, policy comparison is fundamentally constrained by a small feasible sample size (e.g., 10 or 50) due to significant human effort and limited inference throughput of policies. This paper proposes a novel statistical framework for rigorously comparing two policies in the small sample size regime. Prior work in statistical policy comparison relies on batch testing, which requires a fixed, pre-determined number of trials and lacks flexibility in adapting the sample size to the observed evaluation data. Furthermore, extending the test with additional trials risks inducing inadvertent p-hacking, undermining statistical assurances. In contrast, our proposed statistical test is sequential, allowing researchers to decide whether or not to run more trials based on intermediate results. This adaptively tailors the number of trials to the difficulty of the underlying comparison, saving significant time and effort without sacrificing probabilistic correctness. Extensive numerical simulation and real-world robot manipulation experiments show that our test achieves near-optimal stopping, letting researchers stop evaluation and make a decision in a near-minimal number of trials. Specifically, it reduces the number of evaluation trials by up to 37% as compared to state-of-the-art baselines, while preserving the probabilistic correctness and statistical power of the comparison. Moreover, our method is strongest in the most challenging comparison instances (requiring the most evaluation trials); in a multi-task comparison scenario, we save the evaluator more than 200 simulation rollouts.
@@ I Introduction
A particularly important aspect of policy evaluation is policy comparison, where two policies are repeatedly deployed in an environment to assess their relative performance. Policy comparison forms the foundation of robot learning as “an empirical science” [28], enabling researchers to objectively measure the scientific progress of the field. Nevertheless, this setting introduces an additional source of stochasticity due to random outcomes of both the first and the second policies, making the reliability of the comparison more challenging to ensure than evaluation of single policies.
@@
To address these challenges, we propose a novel sequential testing framework named STEP (Sequential Testing for Efficient Policy Comparison) for rigorously comparing performance of imitation learning policies [^1] Although this paper focuses on imitation learning, STEP is naturally applicable to evaluating any types of policies based on binary metrics, including reinforcement learning (RL) policies with sparse 0/1 reward.. Unlike batch testing, our approach allows the number of trials to be varied within a given experimental budget. This critical feature offers two practical advantages. First, the evaluator can stop conducting trials early without sacrificing the probabilistic correctness of the comparison if enough statistical evidence is accumulated quickly in favor of $\pi_{1}$ (or $\pi_{0}$). Second, it reduces the epistemic risk of overconfident (and potentially incorrect) conclusions when $\pi_{0}$ and $\pi_{1}$ are closely matched, since the test will abstain from making a decision if statistical evidence remains low after all the planned trials have been executed. Alternatively, the evaluator may safely append additional trials to the original samples and re-conduct statistical analysis on all the results obtained thus far without inadvertent p-hacking. We demonstrate these advantages through simulation and real-world robot manipulation experiments. Furthermore, extensive numerical experiments show that STEP significantly outperforms state-of-the-art (SOTA) sequential methods, reducing the required number of trials by up to 37% without sacrificing probabilistic correctness. The specific contributions of this paper are as follows:
  • We propose STEP, a novel sequential statistical framework for evaluating relative performance of two policies with tunable probabilistic correctness  [^2] Associated website and code can be found at: https://tri-ml.github.io/step/.
@@ II Preliminaries
Null Hypothesis $\begin{split}\text{Null Hypothesis }\mathbb{H}_{0}:&\hskip 5.69054ptp_{1}~{}\leq~{}p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{0}\\ \text{Alt. Hypothesis }\mathbb{H}_{1}:&\hskip 5.69054ptp_{1}~{}>~{}p_{0}\equiv(p_{0},p_{1})\in\mathcal{H}_{1}.\end{split}$ (3)
@@
Finally, while the preceding development has focused on one-sided testing, the comparison problem naturally admits a bidirectional version allowing decisions for membership in either the alternative set (RejectNull) or the null set (AcceptNull). Note that the decision AcceptNull formally amounts to rejecting the null of a “flipped” version of Eq. 3. In all subsequent discussion we will implicitly utilize the bidirectional test, which will allow for the decision AcceptNull.
@@ IV Methodology
There are several important practical considerations to constructing a near-optimal solution to Eq. 4. We present the concrete challenges first, and then discuss the technical innovations that account for them. Throughout, we will leverage significant mathematical structure in the testing problem. Where insightful or intuitive, this will be explained in situ. See Section X-F in the Supplement for additional details.
@@ IV-C Type-I Error Control
The composite nature of the null hypothesis (which contains all pairs $\mathcal{H}_{0}=\{(p_{0},p_{1})\in[0,1]^{2}\text{ }\rvert\text{ }p_{0}~{}\geq~{}p_{1}\}$) means that the decision-making problem can be thought of (informally) as analogous to distributionally robust control, where the uncertainty is over the particular worst-case $(p_{0},p_{1})\in\mathcal{H}_{0}$. Controlling the Type-I Error in policy comparison then amounts to controlling the Type-I Error uniformly (robustly) for all $h\in\mathcal{H}_{0}$. Suppose that some rejection region for the first $n-1$ steps has been obtained with accumulated risk $\alpha_{n-1}$, and we are interested in bounding the Type-I Error for $n$ by some $\alpha_{n}>\alpha_{n-1}$. Given the notion of stopping and continuation regions introduced in the previous section, this is expressed mathematically as:
@@
The dependence on the (probabilistic) rejection region for $n-1$ is made explicit in Eq. 7, reflecting the internal dynamic structure. For example, if some state $\mathbf{x}_{n-1}=(a,b,n-1)$ is rejected at $n-1$ with a non-zero probability, then the 1-step reachable states (e.g., $(a+1,b,n)$) under the dynamics Eq. 5 are less likely to be feasible at $n$. The presence of many (infinite) elements $h$ in the set of null hypotheses $\mathcal{H}_{0}$ makes verification of Type-I Error control challenging a priori, even for a single particular $n$. However: monotonicity, symmetry, and smoothness properties of this family of testing problems (noted in, for example, [13]) allow for efficient discretization procedures that preserve safety. Specifically, it suffices to consider a discrete set of the “worst-case” nulls ^$\hat{\mathcal{H}}_{0}=\{(p^{(1)},p^{(1)}),(p^{(2)},p^{(2)}),\cdots,(p^{(M)},p^{(M)})\}$, where $10~{}\leq~{}p^{(1)}\cdots~{}\leq~{}p^{(M)}~{}\leq~{}$ 1. Further details of this discretization process is given in Section X-F2 in the Supplement. Thus, Eq. 7 reduces to a finite set of individual inequality constraints.
@@ IV-E Tractable Optimization
Algorithm 1 STEP Decision-Rule Synthesis
  Input: $N_{\text{max}}>0$, risk budget function $f(n)$, type-I error limit $\alpha^{*}\in(0,1)$, number of approximation points $M$
  Initialize: $\zeta_{0}=\emptyset$, $(P_{0})_{ij}=1$ if $(i,j)=(0,0)$ else 00.
  for n $\in\{1,...,N_{\text{max}}\}$ do
     $P_{n}\leftarrow\textbf{Propagate}(P_{n-1},\zeta_{n-1},M)$Eq. 8}
     $\mathbf{w}_{n}\leftarrow\textbf{Opt}(P_{n},f)$Eq. 10}
     $\zeta_{n}\leftarrow\textbf{Compress}(\mathbf{w}_{n})$
  end for
  return  $\zeta=\{\zeta_{1},\dots,\zeta_{N_{\text{max}}}\}$STEP policy}
The optimization problem Eq. 10 is a simple linear program that can be efficiently solved by a standard optimization software; it is a maximization of $\lVert\mathbf{w}_{n}\rVert_{1}$ over the nonnegative orthant, subject to additional linear inequality constraints to control Type-I Error. Because each $P_{n}$ depends on the corresponding $\zeta_{n-1}$, the optimization needs to be performed sequentially for each $n$ in increasing order. Nevertheless, all the computation can be performed offline prior to running the actual policy evaluation. See Algorithm 1 for the decision-rule synthesis procedure. The fact that the optimization problems are sequentially solvable is owed to the isolation of the risk accumulation rate $f(n)$ as a tunable parameter; otherwise, the rejection regions would be generally coupled across $n$ and the optimization would be more challenging. Finally, we note in Algorithm 1 that the representation of the optimal policy can be compressed substantially from the vector $\mathbf{w}_{n}$ to a set of connected sets $\zeta_{n}=\bigg{\{}\mathcal{X}^{\text{Reject Null}}_{n},\mathcal{X}^{\text{Accept Null}}_{n},\mathcal{X}^{\text{Continue}}_{n}\bigg{\}}$ defined by their boundaries, due primarily to the general monotonicity of the optimal regions and the unimodality of the distribution of the test statistic.
@@ V-A Baseline Procedures
The baselines constitute the SOTA sequential analysis methods described in Lai [30] and Lai and Zhang [31] (termed “Lai”) and the Safe, Anytime-Valid Inference (SAVI) method of Turner and Grünwald [50], which is specifically tailored to contingency table (i.e., policy comparison) problems (termed “SAVI”). Similar to STEP, Lai is a valid sequential method under a pre-determined $N_{\max}$ and is asymptotically optimal as $N_{\max}$ tends to infinity. The specific difference between these methods is that Lai makes the asymptotic assumption of normality of the empirical means. This allows for any distribution subject to the central limit theorem (CLT) to be analyzed tractably in the large-data regime, as the resulting optimal stopping PDE amounts to the solution of a reverse heat equation. STEP, by contrast, solves the exact PDE at the fundamental finite-sample resolution of the testing problem. As will be shown, the Lai procedure does best in difficult (i.e., requiring a lot of data) and symmetric (i.e., $p_{0},p_{1}$ each near $0.5$) test settings, as these best match the asymptotic approximation where the CLT approximation is well-matched to the exact distribution of the empirical means. SAVI, by contrast, does not require a fixed $N_{\max}$ a priori and maintains validity for an arbitrarily large $N_{\max}$. In essence, SAVI procedures construct a test statistic subject to stochastic dynamics (the sequential incorporation of data) which possesses a high-probability ‘stability certificate’ under any null data generating process. The ‘stability’ in this case grants time-uniformity (in perpetuity, i.e., not requiring any finite $N_{\text{max}}$); thus, under any null hypothesis, the probability of ever violating the certificate is bounded (by design) to be less than $\alpha$. In exchange for increased generality and robustness, however, SAVI methods inevitably sacrifice sample complexity in cases where a finite upper-bound to $N_{\max}$ is determined by practical time and resource constraints that researchers are subject to. We emphasize that each existing method is widely applicable to practical testing problems beyond robotics, as is STEP. Importantly, STEP is tailored specifically to the small-$N_{\text{max}}$ evaluation regime.
Additionally, an Oracle Sequential Probability Ratio Test (SPRT) [55] is run using the true singleton alternative and associated worst-case singleton null; this method represents a near-optimal procedure for easier problems where feasible methods quickly approach a terminal power close to 1 and still serves as a reasonable benchmark in harder problems. We emphasize that in practical cases, the Oracle method is infeasible to the evaluator; it is included to give a conservative estimate of the optimality gap of each method. Additional information about each baseline is included in Section VI.
@@ V-B Numerical Simulation
The algorithms are first validated on multiple pairs of $(N_{\text{max}},\alpha^{*})$. The ensuing numerical results will use the case $N_{\text{max}}=500,\alpha^{*}=0.05$; similar figures for $N_{\text{max}}=100,\alpha^{*}=0.05$ are included in Section X-B in the Supplement. For each method, we compute for each of the 45 alternatives the following characteristics: (1) Type-I Error – the fraction of associated worst-case null trajectories which have incorrectly rejected the null; (2) Terminal Power – the fraction of alternative trajectories which have correctly rejected the null by step $N_{\text{max}}$; (3) Cumulative Power – a visualization of the fraction of alternative trajectories which have correctly rejected the null by step $t$ for all $t\in\{0,1,...,N_{\text{max}}\}$. Note that another natural evaluation metric, the expected time-to-decision ($\mathbb{E}[n_{\text{stop}}]$), can be derived from the cumulative power as the area between the curve and the constant $y=1$.
@@ V-B2 What are the unique advantages of STEP in comparison problems of varying difficulty?
We now demonstrate the downside of the Lai baseline procedure as compared to our method through the cumulative power of the procedure.
In Fig. 5 (right), we illustrate the cumulative power on a hard evaluation case. This represents a second view of the observation presented in the terminal power analysis: the Lai procedure and our STEP each significantly outperform the SAVI procedure in the small-gap regime.
Conversely, in Fig. 5 (center), we illustrate a difficult setting in which the Lai procedure struggles. In this setting the performance gap is again 10 percentage points, but the distribution is lower-variance and skewed as compared to that of Fig. 5 (right); as such, the Lai procedure cannot effectively adapt and our STEP significantly outperforms in terms of deciding more quickly for the alternative.
@@ V-C How sample efficient is STEP in practical policy comparison settings?
In addition to the numerical validation, we evaluate STEP through two sets of real-world robot evaluation experiments and a simulation experiment. In the first set of experiments, we compared policies with noticeable performance gaps to demonstrate the early-stopping capability of our approach. In the second set, we compared closely-competing policies to characterize necessary sample sizes for statistical validity when policy performance gaps become small. In the simulation experiment, we perform multi-task comparison of two SOTA imitation learning policies. See Section X-A in the Supplement for more details on the hardware experiments.
@@ V-C1 Hardware Evaluation in the Large/Medium-Gap Regime
In FoldRedTowel, we compare two policy checkpoints from a single training run. The baseline policy $\pi_{0}$ was trained for 10000 gradient steps with an AdamW [34] optimizer and achieved the action mean-squared error (MSE) of 1.61E-3 on the validation set. The other policy $\pi_{1}$ continued the training for 70000 additional steps, yielding the validation action MSE of 1.35E-3. To evaluate each policy, five in-distribution (ID) initial conditions were chosen and repeated 10 times each, constituting 50 total trials. ID initial conditions visually resemble the ones that appear in the training dataset. As shown in Table I (rows 1-3), the empirical gap in success rates was 36 percentage points (56% to 92% success), suggesting that $\pi_{0}$ was under-trained. Each sequential method detected a significant difference at level $\alpha^{*}=0.05$ in 19 – 23 trials (apiece). That is, had the algorithms been active during collection, the last 27 – 31 rollouts per policy were unnecessary for confirming the improvement of the later checkpoint over the former. Additionally, note that the Lai method and our STEP sequential procedures were each tuned for an $N_{\text{max}}$ of 50 (row 1), 200 (row 2), and 500 (row 3) rollouts. In the latter two cases, additional rollouts could have been run up to 200 or 500 per policy if the gap was smaller without compromising the validity of the decision.
@@
FoldRedTowel $\pi_{i}$ 0.05 500 50 0.560 0.920 20 23 23 17
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 50 50 0.280 0.800 7 9 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 200 50 0.280 0.800 7 12 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 500 50 0.280 0.800 7 14 14 7
@@
TABLE I: Empirical time-to-correct-decision for all hardware (top) and simulation (bottom) policy comparisons. The comparison type is described first; $\pi_{i}$ is comparing two policies, $\mathcal{D}^{i}_{s_{0},o_{0}}$ compares one policy under possible distribution shift. The utilized Type-I Error $\alpha^{*}$ and $N_{\text{max}}$ describe the constraints applied a priori by the evaluator (we underline to emphasize the change in $N_{\text{max}}$ for rows 1-3 and 4-6; observe that the sensitivity of the stopping times is very small). $N~{}\leq~{}N_{\text{max}}$ represents the amount of data available for the statistical analysis. We report the terminal empirical success rates (after $N$ trials) of each policy in each setting under $\hat{p}_{i}$ (this information is not available to any feasible algorithm). We do not have truth labels on this data; however, in all cases, every method arrived at the same decision, including the Oracle SPRT which has a priori access to privileged information $(\hat{p}_{0},\hat{p}_{1})_{N}$; this decision was Reject Null for all rows except the CarrotOnPlate tasks, which each returned Fail To Decide. We report the stopping times of all methods on the right of the table for every context; in all cases: lower is better. We put in bold any feasible method result that is near-optimal within ten trials (absolute) or 25% (relative) of the SPRT Oracle, which is not implementable by an evaluator. In the Multitask setting, we test $p_{1}>p_{0}$ uniformly across the preceding three tasks. This stopping time is the sum by column of the stopping times for the three tasks. Our method saves the evaluator over 200 trials in uniform certification over these three tasks as compared to either feasible baseline.
In CleanUpSpill, we compare the same policy on two different sets of initial conditions. The task is similar to the one originally presented by Xu et al. [60], which compares a set of ID initial conditions against the out-of-distribution (OOD) initial conditions. The ID set includes 10 initial conditions with a white towel and a short blue mug whereas the OOD set uses 10 with a checkered towel with a tall cyan mug (each initial condition is repeated five times). As shown in Table I (rows 4–6), the empirical gap of 52 percentage points was detected in 7–14 trials by all methods, though they were tuned (where applicable) to an $N_{\text{max}}$ up to thirty to seventy times larger; this demonstrates the significant reduction in sensitivity (from an evaluator’s standpoint) arising from setting $N_{\text{max}}$ versus choosing a batch size $N$. Furthermore, STEP’s efficiency only minimally degrades when $N_{\text{max}}$ is increased from 200 (row5) to 500 (row 6). In this setting, any of the three sequential methods would have prevented the need for nearly 70 of the 100 total rollouts (35 of the 50 batch trials per policy).